OpenStudy (anonymous):

Solve the equation exactly in terms of log. 10^(6-2x) = 3 x= ?

OpenStudy (anonymous):

log 3=6-2x (log 3)-6 = 2x ((log 3)-6)/2) = x

OpenStudy (anonymous):

this is because log = 10^ correct?

OpenStudy (anonymous):

its log base 10 so log X 10

OpenStudy (anonymous):

ah thanks